# From Large to Super-Tiny: End-to-End Optimization for Cost-Efficient LLMs

### v1

@@ Abstract
In recent years, Large Language Models (LLMs) have significantly advanced artificial intelligence by optimizing traditional Natural Language Processing (NLP) pipelines, improving performance and generalization. This has spurred their integration into various systems. Many NLP systems, including ours, employ a "one-stage" pipeline directly incorporating LLMs. While effective, this approach incurs substantial costs and latency due to the need for large model parameters to achieve satisfactory outcomes. This paper introduces a three-stage cost-efficient end-to-end LLM deployment pipeline—including prototyping, knowledge transfer, and model compression—to tackle the cost-performance dilemma in LLM-based frameworks. Our approach yields a super tiny model optimized for cost and performance in online systems, simplifying the system architecture. Initially, by transforming complex tasks into a function call-based LLM-driven pipeline, an optimal performance prototype system is constructed to produce high-quality data as a teacher model. The second stage combine techniques like rejection fine-tuning, reinforcement learning and knowledge distillation to transfer knowledge to a smaller 0.5B student model, delivering effective performance at minimal cost. The final stage applies quantization and pruning to extremely compress model to 0.4B, achieving ultra-low latency and cost. The framework’s modular design and cross-domain capabilities suggest potential applicability in other NLP areas.
[…]
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have made significant strides [1; 2; 3] in enhancing artificial intelligence by addressing shortcomings of traditional pipelines — including manual annotation, suboptimal performance and limited generalization and adaptability. Consequently, LLM-based pipelines have stimulated substantial demand for integration into various Natural Language Processing (NLP) systems. We find that numerous NLP systems, including our own, have directly adopted LLM into the pipeline, often referred to as the "one stage" pipeline. Despite the breakthroughs facilitated by this pipeline, it also incurs substantial cost increases [4] and high inference latency [5]. Within our NLP system, cost-performance is an especially critical metric, a sentiment likely shared by others in the field [6]. These significant costs limit the broader applicability of LLMs and impact their commercial viability. Our analysis points that it is precisely the simplistic training method of the one-stage pipeline, which necessitates larger model parameters to achieve satisfactory results, that leads to significant costs.
@@
In this paper, we introduces three-stage performant and cost-efficient end-to-end LLM deployment pipeline, consisting of prototyping, knowledge transfer, and model compression, to address the pervasive "cost-performance" dilemma in practical deployment of LLM-based frameworks. This approach ultimately yields a super tiny model, which is deployed in our online system, achieving optimal cost-performance metrics while substantially simplifying system architecture. The first stage involves constructing the prototype system, where traditional complex tasks are converted into function call-based LLM-driven workflows [7] through prompt engineering, resulting in an optimal performance prototype compared to traditional pipeline. This prototype serves as a teacher model in the second stage, efficiently generating vast quantities of high-quality data. During the knowledge transfer phase, leveraging the optimal prototype from the first stage, we employ techniques including rejection fine-tuning (RFT), reinforcement learning (RL) [8; 9; 10], and multi-strategy knowledge distillation (KD) targeting the Qwen 2.5 series model [11], effectively transferring knowledge to a 0.5B student model. The derived student model is considerably smaller than the prototype model while achieves nearly comparable performance at minimal cost. In the third stage, we utilize quantization and pruning techniques to further compress the student model developed in the second phase, resulting in a remarkably compact model of 0.4 billion parameters (as shown in Figure 1). Ultimately, this super tiny model is deployed in our online system. This model achieves hundreds-fold compression and extremely low latency and cost in exchange for a very small and acceptable loss in performance. The modular design and superior cross-domain performance of the framework suggest its potential applicability in other NLP domains. The major contributions of this paper are as follows:
  • The first validated three-stage cost-efficient end-to-end LLM deployment pipeline that resolves the "cost-performance" dilemma through integrated technical practice, offering new insights for potentially cross-domain industrial-level solutions.
  • Introduction of an LLM-based generation paradigm replacing BERT classification for sample generation and evaluation, that leverages the benefits of LLM technology in line with current trends and enhances NLP performance across multi-domain scenarios.
  • Extensive experiments demonstrating state-of-the-art results, achieving an 180-fold compression with nearly consistent AR compared to 72B LLMs and improving accuracy by an absolute increase of 13% compared to traditional BERT-based pipelines.
@@ 2 Related Works
Recent works [8; 9; 10] focus on leveraging Reinforcement Learning (RL) to enhance model reasoning capabilities. Notable reasoning models such as DeepSeek-R1 [43], QWQ-32B [44], and OpenAI o1 [45] have emerged, employing RL algorithms like Proximal Policy Optimization (PPO) [46], Group Relative Policy Optimization (GRPO) [47], and others [48; 49; 50]. For general-purpose LLMs, preference optimization algorithms [51; 52; 53; 54] are widely adopted.
@@ 3 Methodology
Recent advancements in LLMs have facilitated significant progress in NLP tasks such as classification and summarization. We harness these capabilities to reconstruct the traditional dialogue system through a cost-efficient end-to-end LLM pipeline, which includes prototyping, knowledge transfer, and model compression, as illustrated in 2(a) and 2(b). Initially, leveraging the exceptional performance of contemporary LLMs, we develop the prototype system. Subsequently, we perform knowledge transfer from large models to smaller models. Finally, we achieve further reduction in model size through additional model compression techniques. Ultimately, our objective is to deliver intelligent user experiences at minimal costs, as demonstrated in 2(c).
@@ 3.1 Prototyping
Traditional dialogue systems typically employ the Intent Classification and Slot Filling (ICSF) framework, often with BERT[57]. However, a single model often fails to handle complex real-world issues, as shown in the Table 1. Therefore, a Domain Classifier (DC) is widely used to first categorize queries into domains, enabling domain-specific ICSF parsing. While the hierarchical DC-ICSF architecture improves accuracy, system complexity rises with feature updates. This results in high inter-module coupling, increasing development and maintenance costs. Moreover, ambiguous domain boundaries make the DC a bottleneck for end-to-end accuracy. This creates a "feature expansion - operational difficulty" feedback loop, hindering dialogue experience optimization.
Modern LLMs demonstrate sufficient capability to address such NLP problems. Particularly in related Tool Use works, LLMs can learn to call functions without training by attaching function descriptions in prompts. In fact, ICSF can essentially be viewed as a function call task, where the model extracts critical information from user queries as function parameters for specific modules to execute. Following this approach, we design a prototype system based on Qwen2.5-72B-Instruct[11]. As illustrated in 2(a), the prototype system accepts user queries and contextual information (including chat history and temporal status information) as input, searches relevant APIs from the API pool, then outputs function names and parameters to call. The traditional system comprises one DC module and multiple ICSF models (based on BERT or regular expressions), each requiring separate maintenance and frequent DC updates. In contrast, the LLM-based system with function call capability handles all processing through a unified model, as demonstrated in Figure 1, which reduces system complexity and improves operational efficiency.
To enhance temporal awareness and mitigate hallucinations, we enrich the input with contextual knowledge via RAG [58], retrieving explanations and extensions of entity-related terms from a curated knowledge base. Instead of adding all callable APIs in the API pool into the LLM prompt, we implement an API retrieval module inspired by Kachuee et al. [15]. The LLM first generates potential API descriptions, then retrieves top-K APIs via vector similarity (using GTE-Qwen2-1.5B [59] as the embedding model and FAISS [60] for indexing). The LLM selects the optimal function and parameters from candidate APIs. If no suitable function exists, the system returns "No appropriate function"; ambiguous queries trigger "Request unclear" responses. For backward compatibility, we design an Adapter to map functions/parameters to legacy domain-intent-slot structures.
@@
To address issues above, we introduce LLM-as-a-judge to evaluate prototype accuracy. We design a Judge based on QwQ-32B[44]: Given context and a set of Function & parameters, the Judge determines whether this combination semantically satisfies the user query. Following Samuylova’s best practices [^1] https://www.evidentlyai.com/llm-guide/llm-as-a-judge, we optimize the Judge using a 1k-expert-annotated dataset. Subsequently, both the LLM and 9 domain experts evaluate 3k test samples under identical conditions. Results show that Judge achieves 93% consistency with expert judgments, with LLM providing logically coherent Chain-of-Thought(CoT) rationales. Based on these findings, we conclude that LLM-as-a-Judge represents the optimal evaluation approach for this task. As demonstrated in Figure 4, we reconstruct the dialogue pipeline from BERT-based DC-ICSF architecture to LLM-based Function Call architecture. Evaluated through LLM-as-a-Judge, the new architecture improves accuracy from 81.43% to 94.74% compared to the legacy system.
@@ 3.2 Knowledge Transfer
Building on our agent-oriented large model design, we propose Supervised Fine-Tuning (SFT) on smaller models using large model-generated data instead of relying solely on in-context learning (ICL) with large models. This approach offers two key benefits: (1) Smaller models can achieve performance comparable to large models when trained on sufficient high-quality data; (2) It significantly boosts inference and optimization efficiency. For a dataset $D$ containing prompt component $x$ and corresponding response component $y$, we define $D(x,y)=\{x_{i},y_{i}\}_{i=1}^{N}$, where $N$ denotes the dataset size. To minimize the Cross Entropy (CE) loss, we first utilize the SFT optimization, and then employ further optimization Rejection sampling Fine-Tuning (RFT). We use LLM judge introduced in Section 3.1 to evaluate training data generated by the LLM prototype, filter out low-quality data, and generate the final training set. The actual training data is formulated as:
@@
Since OpenAI introduced Reinforcement Learning from Human Feedback (RLHF) [61], it has become an essential component for aligning LLMs, as evidenced in major technical reports [62; 63; 64; 11; 65]. DeepSeek-R1 [43] proposed a cost-effective RL training method from scratch which combined rule-based ORM and GRPO [47]. Subsequent works like Logic-RL [8; 9; 10] also successfully applied this efficient approach across various tasks. We adapt this method to allow the model to automatically generate Chain of Thought (CoT) during training, which enhances performance [66] and provides insights into model reasoning for subsequent optimization. Also, we primarily follow Logic-RL [9], refining the reward design for this specific task (see Appendix B) while utilizing the same training dataset as RFT to achieve enhanced model performance.
@@ 3.3 Model Compression
To further compress model size, we design a two-stage pruning pipeline following NVIDIA [67], Llama3.2 [68], and Apple [69]. Offline One-Shot Struct Pruning: Following Minitron[31], we explore two pruning strategies, for Depth Pruning, we compute layer importance scores and remove the $N$ least important layers based on the target pruning ratio, for Width Pruning, we conduct a model architecture search using PPL-based criteria and prune channels according to group-wise importance rankings, with full implementation details in Section I.1. Recovery training: After offline one-shot pruning, we execute a two-phase recovery process: (1) General Instruction Fine-tuning via Knowledge Distillation: Adopting Minitron’s approach, we leverage logits from a larger general instruction teacher model for recovery training. Due to unavailability of the teacher’s original training data, we first perform model correction on our selected IFT dataset to stabilize distillation, followed by top-k logits distillation(Section 3.2); (2) Domain-Specific Task Alignment: The pruned compact LM from phase 1 undergoes further aligning through knowledge distillation, producing a final domain-optimized LLM.
@@
Additionally, quantization significantly reduces GPU memory consumption and enhances inference performance. In this work, we apply post-training quantization to the smaller models obtained through distillation/pruning. Specifically, we implement two quantization approaches: 4-bit weight-only GPTQ algorithm [33] and full 8-bit weight-activation FP8 quantization [41]. The 4-bit weight-only quantization [33] compresses linear weights from FP16 to per-group INT4 via integer quantization, which are dequantized to FP16 during computation. This method achieves notable inference acceleration in memory-bound scenarios with large model parameters and low GPU memory bandwidth. The FP8 quantization [41] quantizes both weights and activations to 8-bit floating-point format. Compared to 8-bit integer quantization [41], FP8 demonstrates superior precision and leverages NVIDIA’s FP8 tensor cores for effective inference acceleration [70].
@@ 4.1 Experimental Setup
We conduct experiments using a large-scale dataset developed by our team, leveraging years of experience and practice in the field. The dataset transforms a real system’s intent recognition and semantic slot framework into an API document specification, creating a function pool with 74 functional modules and 225 parameters. We sample 200,000 data points from a single business scenario as the training set, with corresponding labels synthesized based on the large language model (LLM) prototype outlined in Section 3.1. Additionally, we employ a combination of hierarchical and random sampling to generate 11,922 samples as an in-domain test set, and 7,867 samples from a different business scenario as an out-domain test set, where some callable functions are absent from the training set. Compared to open-source datasets, our custom dataset is more intricate and closely mirrors real-world scenarios (see Table 1).
@@
We select small models (0.5B-7B) from the Qwen2.5 model family [11] for RFT, perform RL on the 3B model, and pruning on the 0.5B model, obtaining Qwen2.5-0.4B-Width and Qwen2.5-0.4B-Depth, both achieved under a 20% pruning ratio. Then, we perform KD on the pruned models and 0.5B model with the 7B RFT-trained teacher. Finally, we apply model quantization using W8A16/W4A16/W8A8 strategies to the KD-trained 0.5B model, detailed in Appendix E. The training process is cost-efficient, utilizing only 8 NVIDIA H20 GPUs. Computing resources are listed in Appendix F.
@@
We evaluate our models with the BERT-Based System, which consists of two models based on BERT-Large (330M)[71]. To evaluate the effectiveness of our method, we invite expert annotators alongside the LLM-as-a-judge-based metric AR (see Equation 6) for performance assessment. We also implement a retrieval module to select the top 10 functions from the function pool, aiming to accelerate both the training and inference phases, as detailed in Appendix D. In real-world scenarios, we deploy models on NVIDIA L20 GPU and perform effectiveness evaluation via TensorRT-LLM[70], with specific configuration details provided in Appendix H. During the comparison between the trained models and a BERT-based system, we adopt queries per second (QPS) as the primary metric for system throughput, response time (RT) for end-to-end latency, and concurrent clients (CC) to quantify high-concurrency performance.

### v2

@@ Abstract
In recent years, Large Language Models (LLMs) have significantly advanced artificial intelligence by optimizing traditional Natural Language Processing (NLP) pipelines, improving performance and generalization. This has spurred their integration into various systems. Many NLP systems, including ours, employ a "one-stage" pipeline directly incorporating LLMs. While effective, this approach incurs substantial costs and latency due to the need for large model parameters to achieve satisfactory outcomes. This paper introduces a three-stage cost-efficient end-to-end LLM deployment pipeline—including prototyping, knowledge transfer, and model compression-to tackle the cost-performance dilemma in LLM-based frameworks. Our approach yields a super tiny model optimized for cost and performance in online systems, simplifying the system architecture. Initially, by transforming complex tasks into a function call-based LLM-driven pipeline, an optimal performance prototype system is constructed to produce high-quality data as a teacher model. The second stage combines techniques like rejection fine-tuning, reinforcement learning, and knowledge distillation to transfer knowledge to a smaller 0.5B student model, delivering effective performance at minimal cost. The final stage applies quantization and pruning to extremely compress models to 0.4B, achieving ultra-low latency and cost. The framework’s modular design and cross-domain capabilities suggest potential applicability in other NLP areas.
* Equal Contribution Corresponding Author
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have made significant strides [1; 2; 3] in enhancing artificial intelligence by addressing shortcomings of traditional pipelines — including manual annotation, suboptimal performance, and limited generalization and adaptability. Consequently, LLM-based pipelines have stimulated substantial demand for integration into various Natural Language Processing (NLP) systems. We find that numerous NLP systems, including our own, have directly adopted LLM into the pipeline, often referred to as the "one stage" pipeline. Despite the breakthroughs facilitated by this pipeline, it also incurs substantial cost increases [4] and high inference latency [5]. Within our NLP system, cost-performance is an especially critical metric, a sentiment likely shared by others in the field [6]. These significant costs limit the broader applicability of LLMs and impact their commercial viability. Our analysis points out that it is precisely the simplistic training method of the one-stage pipeline, which necessitates larger model parameters to achieve satisfactory results, that leads to significant costs.
@@
In this paper, we introduce a three-stage performant and cost-efficient end-to-end LLM deployment pipeline, consisting of prototyping, knowledge transfer, and model compression, to address the pervasive "cost-performance" dilemma in the practical deployment of LLM-based frameworks. This approach ultimately yields a super tiny model, which is deployed in our online system, achieving optimal cost-performance metrics while substantially simplifying system architecture. The first stage involves constructing the prototype system, where traditional complex tasks are converted into function call-based LLM-driven workflows [7] through prompt engineering, resulting in an optimal performance prototype compared to traditional pipeline. This prototype serves as a teacher model in the second stage, efficiently generating vast quantities of high-quality data. During the knowledge transfer phase, leveraging the optimal prototype from the first stage, we employ techniques including rejection fine-tuning (RFT), reinforcement learning (RL) [8; 9; 10], and multi-strategy knowledge distillation (KD) targeting the Qwen 2.5 series model [11], effectively transferring knowledge to a 0.5B student model. The derived student model is considerably smaller than the prototype model while achieving nearly comparable performance at minimal cost. In the third stage, we utilize quantization and pruning techniques to further compress the student model developed in the second phase, resulting in a remarkably compact model of 0.4 billion parameters (as shown in Figure 1). Ultimately, this super tiny model is deployed in our online system. This model achieves hundreds-fold compression and extremely low latency and cost in exchange for a trivial and acceptable loss in performance. The modular design and superior cross-domain performance of the framework suggest its potential applicability in other NLP domains. The major contributions of this paper are as follows:
  • The first validated three-stage cost-efficient end-to-end LLM pipeline with simplifying original system complexity achieves SOTA results and resolves the "cost-performance" dilemma, realizing a 180-fold compression with nearly consistent AR compared to 72B LLMs and up to an absolute 14% AR increase compared to traditional BERT-based pipelines.
  • We propose the first hybrid strategy that significantly enhances the generalization of super tiny models on domain-specific tasks through combining RL and KD. This approach effectively transfers the capabilities of LLMs to smaller models.
  • The proposed framework offers advantages such as low training, inference, and deployment costs, minimal project development expenses with fast cycles, and a strong potential for cross-domain transferability. It provides inspiration for developing standardized cross-domain solutions in both academia and industry, facilitating the widespread LLMs adoption.
@@ 2 Related Works
Recent works [8; 9; 10] focus on leveraging Reinforcement Learning (RL) to enhance model reasoning capabilities. Notable reasoning models such as DeepSeek-R1 [43], QWQ-32B [44], and OpenAI o1 [45] have emerged, employing RL algorithms like Proximal Policy Optimization (PPO) [46], Group Relative Policy Optimization (GRPO) [47], and others [48; 49; 50]. For general-purpose LLMs, preference optimization algorithms [51; 52; 53; 54] are widely adopted. Additionally, self-improving pipelines [55; 56; 57] have been proposed to iteratively refine model performance. We primarily utilize RL to enhance the model’s generalization. Compared to directly distilling a reasoning model, RL results in shorter and more precise Chains-of-Thought (CoT).
@@ 3 Methodology
Recent advancements in LLMs have facilitated significant progress in NLP tasks such as classification and summarization. We harness these capabilities to reconstruct the traditional dialogue system through a cost-efficient end-to-end LLM pipeline, which includes prototyping, knowledge transfer, and model compression, as illustrated in 2(a) and 2(b). Initially, we develop the prototype system by leveraging the exceptional performance of contemporary LLMs. Subsequently, we perform knowledge transfer from large models to smaller models. Finally, we achieve additional reduction in model size through further model compression techniques. Ultimately, our objective is to deliver intelligent user experiences at minimal costs, as demonstrated in 2(c).
@@ 3.1 Prototyping
Traditional dialogue systems usually employ the Intent Classification and Slot Filling (ICSF) framework, often utilizing BERT[60; 61]. However, a single model often struggles to handle complex real-world issues, as shown in the Table 1. Consequently, a Domain Classifier (DC) is frequently used to first categorize queries into domains, enabling domain-specific ICSF parsing. Although the hierarchical DC-ICSF architecture improves accuracy, the system complexity increases with feature updates. This leads to high inter-module coupling, increasing development and maintenance costs. Furthermore, ambiguous domain boundaries turn the DC into a bottleneck for overall end-to-end accuracy, creating a "feature expansion - operational difficulty" feedback loop that hinders dialogue experience optimization.
Modern LLMs exhibit adequate capability to tackle these NLP problems. Particularly in related Tool Use works, LLMs can learn to call functions without additional training by attaching function descriptions in prompts. Essentially, ICSF can be viewed as a function call task, where the model extracts critical information from user queries as function parameters for specific modules to execute. Following this approach, we design a prototype system based on Qwen2.5-72B-Instruct[11]. As illustrated in 2(a), the prototype system accepts user queries and contextual information (including chat history and temporal status information) as input, searches relevant APIs from the API pool, then outputs function names and parameters to call. The traditional system comprises one DC module and multiple ICSF models (based on BERT or regular expressions), each requiring separate maintenance and frequent DC updates. In contrast, the LLM-based system, equipped with function call capability, processes all tasks through a unified model, as illustrated in Figure 1, thereby reducing system complexity and enhancing operational efficiency.
To enhance temporal awareness and mitigate hallucinations, we enrich the input with contextual knowledge via RAG [62], retrieving explanations and extensions of entity-related terms from a curated knowledge base. Instead of adding all callable APIs from the API pool into the LLM prompt, we implement an API retrieval module inspired by [15]. The LLM first generates potential API descriptions, then retrieves top-K APIs via vector similarity (using GTE-Qwen2-1.5B [63] as the embedding model and FAISS [64] for indexing). Then, the LLM selects optimal functions/parameters from candidate APIs. If no suitable function exists, the system returns "No appropriate function"; in the case of ambiguous queries, "Request unclear" are generated. For backward compatibility, we design an Adapter to map functions/parameters to legacy domain-intent-slot structures.
@@
To address issues above, we introduce LLM-as-a-judge to evaluate prototype accuracy. We design a Judge based on QwQ-32B[44]: Given context and a set of Function & parameters, the Judge determines whether this combination semantically satisfies the user query. Following Samuylova’s best practices [^1] https://www.evidentlyai.com/llm-guide/llm-as-a-judge, we optimize the Judge using a 1k-expert-annotated dataset. Subsequently, both the LLM and 9 domain experts evaluate 3k test samples under identical conditions. Results show that Judge can provide logically coherent Chain-of-Thought(CoT) rationales and it achieves 93% consistency with expert judgments. Based on these findings, we conclude that LLM-as-a-Judge is the optimal evaluation approach for this task. As demonstrated in Figure 4, we reconstruct the dialogue pipeline from BERT-based DC-ICSF architecture to LLM-based Function Call architecture. Evaluation through LLM-as-a-Judge demonstrates that the new architecture improves accuracy from 81.43% to 94.74% compared to the original one.
@@ 3.2 Knowledge Transfer
Building on our agent-oriented large model design, we propose Supervised Fine-Tuning (SFT) on smaller models using large model-generated data instead of relying solely on in-context learning (ICL) with large models. This approach offers two key benefits: (1) Smaller models can achieve performance comparable to large models after trained on sufficient high-quality data; (2) It significantly boosts inference and optimization efficiency. For a dataset $D$ containing prompt component $x$ and response component $y$, we define $D(x,y)=\{x_{i},y_{i}\}_{i=1}^{N}$, where $N$ denotes the dataset size. To minimize the Cross Entropy (CE) loss, we first utilize the SFT optimization, and then employ further optimization Rejection sampling Fine-Tuning (RFT). We use LLM judge introduced in Section 3.1 to evaluate training data generated by the LLM prototype, filter out low-quality data, and generate the final training set. The actual training data is formulated as:
@@
Since OpenAI introduced Reinforcement Learning from Human Feedback (RLHF) [65], it has become an essential component for aligning LLMs, as evidenced in major technical reports [66; 67; 68; 11; 69]. DeepSeek-R1 [43] proposes a cost-effective RL training method from scratch which combined rule-based ORM and GRPO [47]. Subsequent works like Logic-RL [8; 9; 10] applies this efficient approach across various tasks, enabling the model to automatically generate Chain of Thought (CoT) during training, thereby enhancing performance [70]. Based on our observation that model generalization degrades after RFT, we attempt to employ RL to improve model performance on out-of-domain datasets. Following Logic-RL [9], we refine the reward design for our specific task (see Appendix B) and conduct extensive attempts to integrate knowledge distillation (see Appendix K).
@@ 3.3 Model Compression
To further compress model, we design two phases of pruning following NVIDIA [71], Llama3.2 [72], and Apple [73]: (1) Offline One-Shot Struct Pruning: Following Minitron[31], we explore two pruning strategies. For Depth Pruning, we compute layer importance scores and remove the $N$ least important layers based on the target pruning ratio. For Width Pruning, we conduct a model architecture search using PPL-based criteria and prune channels according to group-wise importance rankings, with full implementation details in Section I.1; (2) Recovery training: The first step is General Instruction Fine-tuning via Knowledge Distillation. Adopting Minitron’s approach, we leverage logits from a larger general instruction teacher model for recovery training. Due to unavailability of the teacher’s original training data, we first perform model correction on our selected IFT dataset to stabilize distillation, followed by top-k logits distillation(Section 3.2). The second step is Domain-Specific Task Alignment. The pruned compact LM undergoes further aligning through knowledge distillation, producing a final domain-optimized LLM.
@@
Quantization typically reduces GPU memory consumption and enhances inference performance. In this work, we apply post-training quantization to smaller models obtained through distillation/pruning. Specifically, we implement two quantization approaches: 4-bit weight-only GPTQ algorithm [33] and full 8-bit weight-activation FP8 quantization [41]. The 4-bit weight-only quantization [33] compresses linear weights from FP16 to per-group INT4 via integer quantization, which are dequantized to FP16 during computation. This approach achieves notable inference acceleration in memory-bound scenarios with large model parameters and low GPU memory bandwidth. The FP8 quantization [41] quantizes both weights and activations to 8-bit floating-point format. Compared to 8-bit integer quantization [41], FP8 shows a more noticeable improvement in accuracy, and by utilizing NVIDIA’s FP8 tensor cores, FP8 can achieve effective inference acceleration [74].
@@ 4.1 Experimental Setup
In our experiments, we utilize a large-scale dataset built from years of team experience and practice accumulated in the field. We reconstruct the intent classification and semantic slot framework within the real-world system into an API document specification, creating a function pool comprising 74 functional modules and 225 parameters. We sample 200k data points from a single business scenario as the training set and synthesize labels based on the LLM prototype outlined in Section 3.1. A combination of stratified and random sampling yields 11,922 samples from this scenario’s data as the in-domain test set. Additionally, we sample 7,867 data points using the same methodology from another business scenario, without some callable functions in the training set, as the out-of-domain test set. Compared to open-source datasets, our self-constructed dataset is more complex and closely aligned with real-world scenarios (see Table 1).
@@
During the training phases, base models are all from small instruct models (0.5B-7B) of the Qwen2.5 model family [11]. We conduct RFT on them and RL on 7B models. We perform pruning on 0.5B models and obtain Qwen2.5-0.4B-Width and Qwen2.5-0.4B-Depth, both under a 20% pruning ratio. Then, we transfer knowledge from 7B RFT-trained models to pruned and 0.5B models via KD. We also apply KD from 7B RL-trained models to 0.5B models, followed by another RL phase (denoted as KD*+RL). Finally, we use model quantization via W8A16/W4A16/W8A8 strategies on the KD-trained 0.5B model, detailed in Appendix E. The training process is cost-efficient, utilizing only 8 NVIDIA H20 GPUs. Computing resources are listed in Appendix F.
@@
We do evaluation by comparing our models with crowdsourced labelers and the BERT-based system consisting of two models based on BERT-large (330M)[75]. To evaluate the effectiveness of our method, we invite expert annotators alongside the LLM-as-a-judge-based metric AR (see Equation 6) for performance assessment. We also implement a retrieval module to select top 10 functions from the function pool, aiming to accelerate both training and inference phases, as detailed in Appendix D. In real-world scenarios, we deploy models on NVIDIA L20 GPU and perform effectiveness evaluation via TensorRT-LLM[74], with specific configuration details provided in Appendix H. During the comparison between the trained models and the BERT-based system, we adopt queries per second (QPS) as the primary metric for system throughput, response time (RT) for end-to-end latency, and concurrent clients (CC) to quantify high-concurrency performance.